OpenStudy (anonymous):

f(x) = 6 + x When fx is 2

OpenStudy (mathteacher1729):

Does the word "evaluate" appear in your problem?

OpenStudy (anonymous):

It says this . Complete the table for the function. With fx = 2 . So you'd have to find x . Does that help ?

OpenStudy (anonymous):

x=2 f(x)=6 + x f(2)=6 + (2) x= 8

OpenStudy (anonymous):

So after adding 6+2, just stubstract it from the fx(2)?

OpenStudy (anonymous):

Just add. They already give you the value of x. So you change x for 2 and solve

OpenStudy (anonymous):

Mmkay . Thanks . Can you help me on the other one ?